OpenStudy (anonymous):

Solve the compound inequality 9 + 3n 3 or 5n < 30. My answer is n −2 or n < 6 Am I right?

OpenStudy (anonymous):

what is the inequality sign between the 3n and the 3 ?

OpenStudy (anonymous):

oops sorry, its ≥

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

9 + 3n >= 3 3n >= 3 - 9 3n >= - 6 n >= -6/3 n >= - 2 5n < 30 n < 6 it can also be written as n >=-2 < 6

OpenStudy (anonymous):

Thank you so much! :)

OpenStudy (anonymous):

your welcome...you didn't need help, you already had the answer. You did a good job :)